How many liters of water should be added to 20L of a 30% acid solution to make a solution that is 10% acid?
Answer by stanbon(75887)

How many liters of water should be added to 20L of a 30% acid solution to make a solution that is 10% acid?
-----------------------
Equation:
acid + acid = acid
----
0.30*20 + 0*x = 0.10(20+x)
------
Multiply thru by 100 to get:
30*20 + 0 = 10*20 + 10x
10x = 20*20
x = 40L (amt. of water that must be added)
===================================
Cheers,
Stan H.